OpenStudy (lena772):

Is the simplified form of 2square root of 3 + 3square root of 3 rational?

OpenStudy (ranga):

\[2\sqrt{3} + 3\sqrt{3} = 5\sqrt{3}\] sqrt(3) is irrational.

OpenStudy (lena772):

therefore the whole equation is irrational?

OpenStudy (ranga):

Yes, the sum is irrational. But it is not an equation though. Just adding two numbers that happens to be irrational.